OpenStudy (anonymous):

A spring requires a force of 25 N to produce an extension of 2 cm in it. What is the work done in further extending it by 3 cm.

OpenStudy (anonymous):

The answer would be 125/2.. How do I convert that to joules?

OpenStudy (anonymous):

F = -kx k = 25/0.02 = 1250 N/m Now to calculate work we'll need to integrate -kx from .02 to .03 \[w = \int\limits_{0.02}^{0.03} -kx dx\] as long as your distance units are meters then your answer will be in Joules.